# MME-RealWorld: Could Your Multimodal LLM Challenge High-Resolution Real-World Scenarios that are Difficult for Humans?

### v1

@@ 3.1 Problem Definition
Let $I\in\mathbb{R}^{H\times W\times C}$ be a high-resolution image, where $H$, $W$, and $C$ denote the height, width, and number of channels, respectively, and let $Q$ be a textual question pertaining to image $I$. The core task in high-resolution multimodal understanding is to generate an accurate answer $A$ based on the joint input $(I,Q)$. Standard MLLMs, denoted by $\mathcal{M}$, often utilize vision encoders accepting fixed-size inputs. This necessitates downsampling high-resolution images $I$ to $I_{\text{ds}}$, where $I_{\text{ds}}\in\mathbb{R}^{h\times w\times C}$ with $h<H$ and $w<W$. The conventional inference process is typically represented as:
@@
This downsampling operation, however, leads to an unavoidable loss of fine-grained visual information present in the original image $I$. Such loss can significantly impair the model’s performance, particularly when the correct answer hinges on discerning small objects, interpreting fine text, or recognizing subtle visual patterns. Our objective is thus to develop a method enabling $\mathcal{M}$ to effectively utilize the rich detail within the original high-resolution image $I$, thereby producing an improved final answer $A_{\text{final}}$ relative to the baseline $A_{base}$, all without necessitating any retraining of the MLLM.
@@ 3.2 Zoom-Refine: Localized Zoom and Self-Refinement
Following the prediction of the bounding box $B$, the Localized Zoom is executed by extracting a high-resolution crop, $I_{\text{crop}}$, directly from the original, full-resolution image $I$ based on the coordinates in $B$. This operation is denoted as:
@@
[…]
@@ 4.1 Evaluation on MME-Realworld Benchmark
As shown in Table 1, Zoom-Refine significantly bolsters the reasoning capabilities of MLLMs on the MME-Realworld benchmark. When applied to InternVL3-78B, Zoom-Refine elevates the weighted average accuracy (Avg) by a substantial $3.60\%$, and the unweighted average accuracy(Avg-C) by an even more impressive $4.39\%$. Similar substantial gains are observed with InternVL2.5-78B, with improvements of $4.37\%$ and $5.07\%$ for Avg and Avg-C respectively. This demonstrates that Zoom-Refine’s core mechanism of providing more granular visual information directly translates to an enhanced ability to perform complex logical deductions. For instance, the marked improvement in Diagram and Table (DT) reasoning (e.g., +$9.60\%$ for InternVL3-78B) suggests that the refined visual input allows the model to better understand structural relationships and extract precise data points crucial for analytical reasoning. Similarly, improvements in other tasks indicate a better grasp of spatial relationships and object interactions, which are fundamental to reasoning in dynamic scenes. The uplift across diverse reasoning subtasks underscores that a more faithful and detailed visual understanding, as facilitated by Zoom-Refine, is a critical precursor to robust multimodal reasoning.
@@ 4.2 Evaluation on HR-Bench Benchmark
On HR-Bench 8K, InternVL3-78B enhanced with Zoom-Refine demonstrates a marked increase in average accuracy, climbing by $5.0\%$ from its baseline. Similarly, InternVL2.5-78B sees $2.9\%$ gain. On the 4K tasks, these models achieved respective gains of $4.2\%$ and $2.1\%$ These results compellingly demonstrate Zoom-Refine’s proficiency in handling the unique challenges posed by extremely high-resolution images, where conventional downsampling or naive patch strategies often lead to critical information loss. The consistent uplift across both FSP, which focuses on discerning attributes of individual small objects, and FCP, which requires comparing details across multiple instances or regions, underscores the comprehensive nature of the improvements. Zoom-Refine empowers MLLMs to effectively navigate and interpret many visual data in such images, leading to more accurate perception of minute details and their contextual relationships.
@@ 4.4 Analysis of Model Scales
To assess Zoom-Refine’s scalability, we applied it to InternVL3 models of different sizes on the MME-Realworld reasoning tasks, with detailed improvements reported in Table 5. Zoom-Refine significantly delivers performance enhancements across the entire spectrum of model sizes. Notably, uplifts are also registered for its smaller models: for $+2.42\%$ the 14B model, $+2.71\%$ for the 8B model, and an impressive $+3.10\%$ for the compact 2B model.

### v2

@@ 3.1 Problem Definition
Let $I\in\mathbb{R}^{H\times W\times C}$ be a high-resolution image, where $H$, $W$, and $C$ denote the height, width, and number of channels, respectively, and let $Q$ be a textual question pertaining to image II. The core task in high-resolution multimodal understanding is to generate an accurate answer AA based on the joint input $(I,Q)$. Standard MLLMs, denoted by $\mathcal{M}$, often utilize vision encoders accepting fixed-size inputs. This necessitates downsampling high-resolution images II to $I_{\text{ds}}$, where $I_{\text{ds}}\in\mathbb{R}^{h\times w\times C}$ with $h<H$ and $w<W$. The conventional inference process is typically represented as:
@@
This downsampling operation, however, leads to an unavoidable loss of fine-grained visual information present in the original image II. Such loss can significantly impair the model’s performance, particularly when the correct answer hinges on discerning small objects, interpreting fine text, or recognizing subtle visual patterns. Our objective is thus to develop a method enabling $\mathcal{M}$ to effectively utilize the rich detail within the original high-resolution image II, thereby producing an improved final answer $A_{\text{final}}$ relative to the baseline $A_{base}$, all without necessitating any retraining of the MLLM.
@@ 3.2 Zoom-Refine: Localized Zoom and Self-Refinement
Following the prediction of the bounding box $B$, the Localized Zoom is executed by extracting a high-resolution crop, $I_{\text{crop}}$, directly from the original, full-resolution image II based on the coordinates in $B$. This operation is denoted as:
@@
Thus, the Zoom-Refine framework transforms the standard single-pass inference into a more sophisticated, two-step reasoning procedure, $\text{Zoom-Refine}(I,Q,\mathcal{M})$, adeptly utilizing the MLLM’s own capabilities to dynamically focus on pertinent high-resolution details and refine its overall understanding.
@@ 4.1 Evaluation on MME-Realworld Benchmark
As shown in Table 1, Zoom-Refine significantly bolsters the reasoning capabilities of MLLMs on the MME-Realworld benchmark. When applied to InternVL3-78B, Zoom-Refine elevates the weighted average accuracy (Avg) by a substantial $3.60\%$, and the unweighted average accuracy(Avg-C) by an even more impressive $4.39\%$. Similar substantial gains are observed with InternVL2.5-78B, with improvements of $4.37\%$ and $5.07\%$ for Avg and Avg-C respectively. This demonstrates that Zoom-Refine’s core mechanism of providing more granular visual information directly translates to an enhanced ability to perform complex logical deductions. For instance, the marked improvement in Diagram and Table (DT) reasoning (e.g., $9.60\%$ for InternVL3-78B) suggests that the refined visual input allows the model to better understand structural relationships and extract precise data points crucial for analytical reasoning. Similarly, improvements in other tasks indicate a better grasp of spatial relationships and object interactions, which are fundamental to reasoning in dynamic scenes. The uplift across diverse reasoning subtasks underscores that a more faithful and detailed visual understanding, as facilitated by Zoom-Refine, is a critical precursor to robust multimodal reasoning.
@@ 4.2 Evaluation on HR-Bench Benchmark
On HR-Bench 8K, InternVL3-78B enhanced with Zoom-Refine demonstrates a marked increase in average accuracy, climbing by $5.0\%$ from its baseline. Similarly, InternVL2.5-78B sees $2.9\%$ gain. On the 4K tasks, these models achieved respective gains of $4.2\%$ and $2.1\%$ . These results compellingly demonstrate Zoom-Refine’s proficiency in handling the unique challenges posed by extremely high-resolution images, where conventional downsampling or naive patch strategies often lead to critical information loss. The consistent uplift across both FSP, which focuses on discerning attributes of individual small objects, and FCP, which requires comparing details across multiple instances or regions, underscores the comprehensive nature of the improvements. Zoom-Refine empowers MLLMs to effectively navigate and interpret many visual data in such images, leading to more accurate perception of minute details and their contextual relationships.
@@ 4.4 Analysis of Model Scales
To assess Zoom-Refine’s scalability, we applied it to InternVL3 models of different sizes on the MME-Realworld reasoning tasks, with detailed improvements reported in Table 5. Zoom-Refine significantly delivers performance enhancements across the entire spectrum of model sizes. Notably, uplifts are also registered for its smaller models: for $2.42\%$ the 14B model, $2.71\%$ for the 8B model, and an impressive $3.10\%$ for the compact 2B model.